OpenStudy (anonymous):

10^x=12 what does x equal?

OpenStudy (anonymous):

show work please.

OpenStudy (anonymous):

\[x \log 10 = \log 12\] \[x = \log 12 \]

OpenStudy (anonymous):

\[\log_{10}\]

OpenStudy (anonymous):

10^x = 12; taking Logarithms to base 10 on both sides we get; log(10^x) = log(12) => x = log(12).

OpenStudy (anonymous):

use properties of logs log base 10 of 12 =x

OpenStudy (anonymous):

because log base a of x=N equals a^N=x http://www.mathwords.com/l/logarithm_rules.htm

OpenStudy (anonymous):

thank you.